OpenStudy (anonymous):

Which property is demonstrated by this equation? 2 + 3 = 3 + 2

OpenStudy (anonymous):

The Distributive Property a(b + c) = ab + ac 2(3 + 4) = 2×3 + 2×4

OpenStudy (mertsj):

The property that says the order of the addends doesn't matter.

OpenStudy (anonymous):

Associative Property (Addition) a + (b + c) = (a + b) + c 2 + (3 + 4) = (2 + 3) + 4

OpenStudy (anonymous):

Associative Property (Multiplication) a(bc) = (ab)c 2(3×4) = (2×3)4

OpenStudy (anonymous):

http://www.purplemath.com/modules/numbprop.htm